### v4

# Iola Walker: A Mobile Footfall Detection System for Music Composition

Abstract
This outing is part of a larger music technology research project. The objective is to find a method for materially enhancing music using hardware and software. There is a strong likelihood that there exists a new medium for experiencing music via a wearable device that ordinary listeners prefer over the current state of the art. If such a medium is discovered, it is a step towards altruistic, prosocial reform in the music industry. A new playback system infrastructure has a chance to soothe some of the societal problems tied to the larger entertainment industry ecosystem. Artifacts for this paper are available here: https://github.com/willbjames/iolawalker
I Introduction: How the new Music Composition and Listening System Works
The project is dubbed "iola walker" in reference to a common polyrhythm, the hemiola. To create an interactive listening experience for the wearer, musicians record songs with a walking listener in mind, imagining a range of walking paces while playing rhythmic parts. The same passages in a musical composition might be recorded multiple times with a variety of underlying pulses. In the associated interactive listening experience, a listener dons a foot-mounted inertial measurement unit and goes for a walk. The Iola Walker mobile app detects their footfalls in real time from the IMU over bluetooth.
Novel music composition gestures can be derived and implemented using this system. Perhaps a playback system could select the version of a recorded musical passage with underlying polyrhythms or "pulse" closest to the listener’s walking pace. Perhaps the crash at the end of a drum fill could be timed to line up precisely with a footstep. Perhaps the next passage of a song could be determined according to a detected change in walking speed that may have been induced by the previous section.
Information about predicted future footfalls is passed from the footfall detection process to the audio playback code running in a render process. This design allows for relevant signals, such as footfalls, to be incorporated into music with zero latency.
This paper documents the process of training a model to detect a walking listener’s footfalls from real-time signal data, as well as exporting that model to work with low latency on an android application. The model was trained on signals from an Mbient Labs foot-mounted IMU [1] collecting data at 200 Hertz, with the ground truth for footfalls annotated by pressing the volume up button on a BLU JL8 android device [2] when the listener’s foot hits the ground. To collect training data, I walked around my neighborhood while clicking the volume up button on an android mobile phone each time my foot hit the ground. I tried several models, and found that an LSTM attained the best performance.
II Related Works: The State of The Art
The signal from a person’s gait has been identified by other teams as a strong candidate for music technology research. But prior research on gait-modulated music ended up focusing on what the wearer was doing more than the music being played.
The Djogger research project was the first to explore this technology in order to help enhance exercise in 2006 [3]. Then several startups in the 2010s including Weav and Rock My Run, created a system for running workouts by changing the speed of songs played to match running pace [4, 5]. Most recently, in the 2020s, Medrhythms, in a partnership with Universal Music Group, has obtained FDA approval for its InTandem music therapy technology to be prescribed for people recovering from a stroke. [6, 7].
This project is different because the focus is not the wearer’s body motions. This project is decidedly not intended to develop a workout aid. It is a method for music composers to augment their art.
Our method to detect the footfall signal so that it is usable in real time is similar to prior LTSM methods documented as being highly accurate [8, 9, 10]. Our training and inference implementation has several unique features stemming from the intention to use it for music. The way to train a model to output a footfall control signal for an audio rendering process is documented in this work.
III Mobile Data Acquisition Pipeline
The shoe-mounted MbientLab IMU streams 200 Hz tri-axial acceleration to the phone via Bluetooth LE. An Android service wraps the sensor API and forwards raw samples to SensorDataManager, whose responsibilities are:
[…]
  • Background thread that drains the queue, batches samples, and flushes to a $64\text{\,}\mathrm{kB}$ buffered CSV writer every $1\text{\,}\mathrm{s}$ or 1000 samples.
  • Footfall annotation—each volume-up press inserts a high-priority event with the tag FOOTFALL.
  • Session management—files are timestamped and recycled automatically to prevent data loss during long recordings.
The resulting CSVs constitute the train200hz.csv and test200hz.csv datasets used in the Python pipeline below.
IV The path to Conv–LSTM
Before arriving at our final Conv–LSTM architecture, we evaluated two simpler approaches, a 1D convolutional neural network and a logistic regression classifier, using the same sliding-window dataset described in Section 2. Although both baselines trained successfully, only the LSTM was accurate enough for real-time inference.
IV-A 1D Convolutional Neural Network
The CNN pipeline (implemented in cnn_pipeline.py) proceeds as follows:
  1. 1.
    Feature extraction: Read raw accelerometer samples ($x,y,z$), compute the magnitude channel $\sqrt{x^{2}+y^{2}+z^{2}}$, and standard-scale all four channels.
  2. 2.
    Windowing: Slice into windows of 200 samples (1 s at 200 Hz) with a stride of 50 samples (75% overlap).
  3. 3.
    Model definition: A three-stage 1D CNN:
    • Two convolution + batch-norm + ReLU blocks with max-pooling.
    • A third convolution with adaptive pooling to collapse the time dimension.
    • Two fully connected layers (64 $\rightarrow$ 64 $\rightarrow$ 1) producing a logit per window.
  4. 4.
    Training: 50 epochs of AdamW (LR = 1e-3), with a ReduceLROnPlateau scheduler and an 80/20 train/validation split.
This model achieved moderate detection accuracy but exhibited limited sensitivity to fine-grained timing.
IV-B Logistic Regression
We also tried a purely linear model (in train_logistic.py):
  1. 1.
    Feature extraction: As above, plus flatten each 50-sample window into a single vector of length $50\times 4$.
  2. 2.
    Model definition: A single linear layer mapping the flattened input to one logit.
  3. 3.
    Training: 100 epochs of AdamW (LR = 1e-3), weighted sampling to correct class imbalance, and a ReduceLROnPlateau scheduler.
While fast to train, logistic regression lacked the representational power necessary for time series data of footfalls, yielding the lowest F1 score.
IV-C Comparative Results
When evaluated on the held-out test set:
$\text{F}_{1}(\text{logistic})<\text{F}_{1}(\text{CNN})<\text{F}_{1}(\text{Conv--LSTM})$
V Conv–LSTM Data Preparation and Script Configuration
This section describes how we load the raw footfall data into our Python pipeline, configure key parameters, label and preprocess the data, and assemble it into the fixed-length windows used for model training.
V-A Dataset Description
We collect two CSV files from the Android app: train200hz.csv (112 655 samples) and test200hz.csv (43 214 samples). Each row contains:
  • Three accelerometer channels: $x,y,z$.
  • A timestamp in milliseconds.
  • An eventType field marking either sample or footfall.
The data are sampled at 200 Hz, giving a new row every 5 ms.
V-B Entry-Point Script Overview
All downstream processing is orchestrated by train_and_test_refined_v2.py, which runs in three sequential stages:
  1. 1.
    Train: fit the Conv–LSTM model with dual (per-sample and per-window) supervision.
  2. 2.
    Export: serialize the best-performing weights to TorchScript and write metadata (window size, stride, threshold, scaler statistics) to JSON.
  3. 3.
    Test: reload the TorchScript model, run a tolerance-sweep evaluation on test200hz.csv, and report precision/recall/F1 at various timing tolerances.
By default, simply invoking:
python train_and_test_refined_v2.py
executes all three stages in order. Command-line flags (e.g. --no-test) can skip individual phases.
V-C Configuration Constants
At the top of the script are defined all hyper-parameters and file paths in one place, facilitating quick experimentation:
Listing 1: Excerpt: configuration block
# Window and stride define how many consecutive samples form one input,
# and how far the window moves each step.
WINDOW_SIZE = 600 # 3 s @ 200 Hz
STRIDE = 150 # 75% overlap
# Training parameters
BATCH_SIZE = 64 # number of windows per batch
LR = 1e-3 # initial learning rate for AdamW
EPOCHS = 50 # number of full passes through the dataset
POS_WINDOW_OVERSAMPLE = 10.0 # oversampling factor for positive windows
# File paths
TRAIN_CSV = "train200hz.csv"
TEST_CSV = "test200hz.csv"
MODEL_OUT = "best_model.pth"
TS_MODEL_OUT = "best_model_ts.pt"
METADATA_OUT = "metadata.json"
Each constant is referenced throughout the code, making it easy to adjust window length, overlap, or training settings without digging into function bodies.
V-D CSV Parsing and Label Assignment
We implement a custom SeqFootfallDataset class to read and label the data:
Listing 2: SeqFootallDataset constructor
class SeqFootfallDataset(Dataset):
def __init__(self, csv_path):
# 1) Read raw CSV, allowing comma or tab delimiters
[…]
# 2) Extract timestamps of the ’footfall’ events
events = df[df.eventType=="footfall"].timestamp.values
# 3) Keep only the continuous ’sample’ rows
samples = df[df.eventType=="sample"].reset_index(drop=True)
# 4) Initialize a target column of zeros
samples["target"] = 0
# 5) For each event time, find the nearest sample index and mark it
for t in events:
idx = (samples.timestamp - t).abs().idxmin()
samples.at[idx, "target"] = 1
# Store the labeled DataFrame for feature extraction
self.samples = samples
This process converts the sparse list of button-press timestamps into a dense binary label for every sample row, which the per-sample head requires.
V-E Feature Engineering
From each row in self.samples we compute five input features:
$x,\;y,\;z,\;\lVert\mathbf{a}\rVert=\sqrt{x^{2}+y^{2}+z^{2}},\;\Delta t=\text{timestamp}_{i}-\text{timestamp}_{i-1}.$
We assemble these into an $(N,5)$ array (where $N$ is the number of samples), then apply a StandardScaler:
  1. 1.
    scaler.fit(training_features) learns mean and variance on the training set only.
  2. 2.
    training_features = scaler.transform(training_features) normalizes each feature to zero mean and unit variance.
  3. 3.
    The same scaler is reused to transform validation and test features, ensuring consistent scaling.
We serialize the fitted scaler to disk (scaler.pkl) so that the mobile inference pipeline applies the identical transform.
V-F Sliding-Window Construction
With scaled features in hand, we generate overlapping windows of length WINDOW_SIZE and stride STRIDE. Internally:
Listing 3: Window slicing in SeqFootfallDataset
self.windows = []
for start in range(0, len(self.samples)-WINDOW_SIZE+1, STRIDE):
# Extract one window of raw features: shape (WINDOW_SIZE, 5)
feats = self.samples.iloc[start:start+WINDOW_SIZE][feat_cols].values
# Extract per-sample labels: shape (WINDOW_SIZE,)
seq_labels = self.samples.iloc[start:start+WINDOW_SIZE].target.values
# Compute a single window-level label: 1 if any target==1, else 0
glob_label = int(seq_labels.max())
self.windows.append((feats, seq_labels, glob_label))
After this loop, $\texttt{len(self.windows)}=\lfloor(N-W)/S\rfloor+1$ windows. The dataset’s __getitem__ returns:
  • $\texttt{feats}\in\mathbb{R}^{\text{WINDOW\_SIZE}\times 5}$,
  • $\texttt{seq\_labels}\in\{0,1\}^{\text{WINDOW\_SIZE}}$,
  • $\texttt{glob\_label}\in\{0,1\}$.
These triples feed directly into the PyTorch DataLoader, which, with our WeightedRandomSampler, produces balanced batches for training.
With data now loaded, labeled, scaled, and windowed, the pipeline proceeds to the training loop described in Section 5.
VI Conv–LSTM Model Definition
In this section we break down each component of the ConvLSTMGlobal architecture in detail.
VI-A Input Representation
Each input to the network is a window of $W=600$ consecutive sensor readings, where each reading has $F=5$ features:
$(x,y,z,\sqrt{x^{2}+y^{2}+z^{2}},\Delta t).$
These are arranged as a tensor of shape $\bigl{(}B,\,W,\,F\bigr{)}$ for a batch of size $B$. We transpose this to (B,F,W)(B,\,F,\,W) before the 1D convolutions.
VI-B 1D Convolutional Layers
The first two layers are 1D convolutions:
  • nn.Conv1d(in_ch, 16, kernel_size=3, padding=1) – This applies 16 different filters of width 3 across the time axis, learning small patterns (e.g. rapid spikes) in the feature channels.
  • nn.ReLU() – A nonlinear activation that zeros out negative values, helping the network model complex relationships.
  • nn.Conv1d(16, in_ch, kernel_size=3, padding=1) – Maps the 16 filtered channels back to the original feature dimension ($F=5$), preparing data for the recurrent stage.
  • nn.ReLU() – Another nonlinearity to enhance representational power.
These convolutions act like learnable sliding windows that detect local temporal patterns in the accelerometer signal.
VI-C Bidirectional LSTM
Next, we feed the convolved features into a bidirectional Long Short-Term Memory (Bi-LSTM):
self.lstm = $n$.LSTM(
input_size=in_ch, # number of feature channels (5)
hidden_size=hid, # size of LSTM hidden state (e.g. 128)
num_layers=2, # two stacked LSTM layers
batch_first=True, # input shape is (B, W, F)
bidirectional=True, # process sequence forwards and backwards
dropout=0.3 # dropout between layers for regularization
  • LSTM cells maintain an internal memory that can capture how sensor readings evolve over time, helping the model recognize the characteristic pattern of a footfall spread over dozens of timesteps.
  • Bidirectional: Each window is processed twice—once from start to end and once in reverse—so the network can use both past and future context when making predictions at each time step.
  • $\texttt{hidden\_size}=hid$ controls the capacity of this memory (the larger, the more complex patterns it can learn, at the cost of more parameters).
The output of the Bi-LSTM at each time step is a vector of length $2\times hid$, combining forward and backward information.
VI-D Linear Heads for Prediction
After the LSTM, we attach two small "heads" that turn the LSTM outputs into actual footfall predictions:
self.seq_head = $n$.Linear(hid*2, 1) # per-timestep prediction
self.global_head = $n$.Linear(hid*2, 1) # whole-window prediction
  • Sequence head (seq_head): Takes the $2hid$-dimensional vector at each of the $W$ timesteps and maps it to a single real number (a "logit") per timestep. After applying a sigmoid, this logit becomes the probability that a footfall occurs exactly at that time index.
  • Global head (global_head): Looks only at the final LSTM output (which summarizes the entire window) and maps it to one logit. After sigmoid, this gives the probability that at least one footfall occurs anywhere in the 600-sample window.
VI-E Why Dual Heads?
  • The sequence head enforces fine-grained timing accuracy, teaching the network exactly which samples correspond to footfalls.
  • The global head provides a coarse binary supervision ("is there any footfall in this window?"), which helps stabilize training when exact timing labels are noisy.
  • Sharing the convolutional and LSTM backbone between both tasks lets the model learn features useful for both detailed and coarse detection, while the small linear heads keep the total parameter count low.
By combining local pattern detection (convolutions), temporal memory (LSTM), and dual objectives (per-sample and per-window), ConvLSTMGlobal balances accuracy with efficiency, making it well-suited for real-time footfall detection on limited hardware.
VII Training Pipeline
Figure 1 illustrates the end-to-end training process. Each block represents one logical stage:
Load &scale CSVSlidewindowsTrain/ValsplitWeightedsamplerEpoch loop(E=50)ThresholdsweepTorchScriptexport
Figure 1: End-to-end Python training flow.
Stage descriptions:
  • Load & scale CSV: Read raw accelerometer data and event annotations from train200hz.csv, then apply a StandardScaler to normalize each feature channel ($x,y,z,\lVert a\rVert,\Delta t$).
  • Slide windows: Segment the continuous, scaled time-series into overlapping windows of length $W$ samples with stride $S$, producing input tensors of shape $(B,W,F)$.
  • Train/Val split: Partition entire walking sessions (files) into training (80%) and validation (20%) sets, stratified so that windows containing at least one footfall are evenly represented.
  • Weighted sampler: Construct a WeightedRandomSampler whose weights boost the probability of sampling windows with positive labels by a factor POS_WINDOW_OVERSAMPLE, mitigating class imbalance.
  • Epoch loop (E=50): For each of the 50 epochs:
    • Iterate over DataLoader batches, each providing:
        inputs $\in\mathbb{R}^{B\times W\times F}$: feature windows.
        seq_targets $\in\{0,1\}^{B\times W}$: per-timestep footfall labels.
        glob_targets $\in\{0,1\}^{B}$: window-level occupancy labels.
    • Forward through ConvLSTMGlobal: yields $\texttt{seq\_logits}\in\mathbb{R}^{B\times W}$ and $\texttt{glob\_logits}\in\mathbb{R}^{B}$.
    • Compute two BCEWithLogitsLoss terms with analytically set pos_weight equal to $\frac{\#\text{neg}}{\#\text{pos}}$ for sequence and global heads, then sum:
      $\mathcal{L}$ $=\frac{1}{B\,W}\sum\mathrm{BCE}(\texttt{seq\_logits},\texttt{seq\_targets})$
      $\quad+\frac{1}{B}\sum\mathrm{BCE}(\texttt{glob\_logits},\texttt{glob\_targets}).$
    • Backward pass: loss.backward(), then clip_grad_norm_(model.parameters(),5.0) to cap gradient norm at 5.
    • Optimizer step: use AdamW (LR = 1e-3, weight decay = 1e-2), then optimizer.zero_grad().
    • Accumulate running losses and any interim metrics (e.g., batch-level precision/recall) for logging.
  • Threshold sweep: After each epoch, disable gradients and run the model over the entire validation set to collect logits. For thresholds $\tau\in\{0.01,0.02,\dots,0.90\}$, compute F1 on binarized $\texttt{seq\_logits}>\tau$. Select the $\tau^{\star}$ maximizing F1.
  • TorchScript export: If the new F1($\tau^{\star}$) exceeds the previous best:
    • Save model.state_dict() as the checkpoint file.
    • Trace the model to TorchScript on CPU using a zero tensor [paszke2019] of shape $(1,W,F)$ and save the serialized .pt file.
    • Write a JSON containing $\{W,S,\tau^{\star},\text{scaler\_mean},\text{scaler\_var}\}$ for mobile inference.
Learning-Rate Scheduling & Logging
  • Attach ReduceLROnPlateau to the optimizer (factor=0.5, patience=5, min_lr=1e-6). After each epoch, call scheduler.step(val_f1).
  • Log to TensorBoard or CSV each epoch:
    • Sequence loss $\mathcal{L}_{\text{seq}}$, global loss $\mathcal{L}_{\text{glob}}$, and total $\mathcal{L}$.
    • Current learning rate.
    • Validation F1 and chosen threshold $\tau^{\star}$.
Invocation
Run the full pipeline via:
python train_and_test_refined_v2.py
Command-line flags such as --no-test or --export-only allow isolating stages for rapid iteration.
VIII Model Export and Metadata Packaging
The best state dict is traced into TorchScript; the scaler and threshold are stored in metadata_refined_v2.json.
Listing 4: TorchScript export
cpu_m = ConvLSTMGlobal().cpu()
cpu_m.load_state_dict(torch.load(MODEL_OUT, map_location="cpu"))
ts = torch.jit.trace(cpu_m, torch.randn(1, WINDOW_SIZE, 5))
ts.save(TS_MODEL_OUT)
IX Evaluation: Tolerance Sweep Tester
The tester rescales test data, runs sliding-window inference, reconstructs predicted indices, and reports precision/recall across tolerances of 0–50 samples (Listing 5).
Listing 5: Tolerance-sweep evaluation loop (excerpt)
for tol in range(0, 51, 5):
tp = fn = fp = 0
f1 = 2*prec*rec/(prec+rec) if prec+rec>0 else 0
results[tol] = {"f1": round(f1, 4)}
X Results Summary
X-A Validation Curve
Table I lists per-epoch validation F1 and threshold.
TABLE I: Validation F1 vs. epoch (best in bold)
Epoch F1 Threshold
1 0.061 0.52
5 0.097 0.78
10 0.107 0.86
15 0.129 0.88
20 0.176 0.89
25 0.226 0.90
30 0.230 0.90
35 0.251 0.90
40 0.252 0.90
47 0.259 0.90
50 0.259 0.90
X-B Test-Set Tolerance Sweep
TABLE II: Test-set metrics vs. tolerance
Tol TP FN FP Prec F1
0 204 74 1195 0.146 0.243
5 255 23 1144 0.182 0.304
10 271 7 1128 0.194 0.323
15 273 5 1126 0.195 0.326
20–50 274 4 1125 0.196 0.327
Allowing a practical timing tolerance (±75 ms) yields a usable precision–recall trade-off. We’ll need timing to be more precise eventually, but for a scenario where we’re choosing passages of music to play in the scale of a few seconds, it’s good enough for now.
XI Android Real-Time Inference Pipeline
The Android application embeds the trained Conv–LSTM model as a TorchScript asset and runs all inference on a dedicated high-priority thread, ensuring minimal end-to-end latency. The core logic lives in the InferenceService class, which performs the following functions in real time:
XI-A Model and Metadata Initialization
Upon startup, InferenceService.initialize() executes on the "InferenceThread" (priority URGENT_AUDIO):
  • Load TorchScript model: Copies refined_v2_ts.pt from the APK’s assets to internal storage and calls Module.load().
  • Parse metadata JSON: Reads metadata_refined_v2.json, extracting
    • windowSize (number of samples per inference window),
    • threshold (offline-tuned global probability cutoff),
    • scaler.mean and scaler.scale arrays for on-device normalization.
  • Buffer pre-allocation: Allocates a flat FloatArray of length $\texttt{windowSize}\times 5$ and a longArrayOf(1,windowSize,5) for the input tensor shape, avoiding any further heap allocations during inference.
  • Audio preparation: Builds a SoundPool with SONIFICATION attributes, preloads a "ding" sample (if available), and falls back to ToneGenerator to guarantee sub-5 ms playback latency.
XI-B Streaming Feature Buffer
As accelerometer data arrives (three floats plus a timestamp), processAccelerometerData(x,y,z,t):
  1. 1.
    Wraps each sample in a small dataWindow deque entry, computing $\Delta t$ from the previous timestamp.
  2. 2.
    Maintains at most windowSize entries, dropping the oldest sample when the buffer is full.
  3. 3.
    Increments a counter and, once inferenceStride new samples have accumulated, posts runModelInference() to the handler thread.
XI-C Model Inference
Inside runModelInference(), executed off the UI thread:
  • Flatten & normalize: Copies the deque into the pre-allocated buffer in $1\times W\times 5]$ layout, then applies $x^{\prime}=\frac{x-\text{mean}}{\text{scale}}$ per channel.
  • TorchScript forward: Constructs an input Tensor and calls module.forward(...), receiving a tuple of:
    • seqLogits (length-$W$ array of per-timestep logits),
    • globalLogit (single logit for the entire window).
  • Sigmoid conversion: Applies $\sigma(z)=1/(1+e^{-z})$ to obtain the per-sample probability $p_{t}$ and window probability $P_{\mathrm{win}}$.
  • Logging: Emits a verbose log of $(p_{t},P_{\mathrm{win}})$ and timestamp, visible via Android’s Logcat for profiling and debugging.
XI-D Decision Logic and Smoothing
Footfall events are declared only when:
  • $p_{t}$ exceeds the sampleThreshold (default 0.85) and $P_{\mathrm{win}}$ exceeds the windowThreshold (default 0.45).
  • A sliding buffer of the last hitWindow boolean decisions contains at least requiredHits positives, implementing an NN-of-$M$ majority filter to suppress isolated spikes.
  • The refractory period of minIntervalMs (e.g. 300 ms) has elapsed since the last confirmed detection.
This multi-stage gating dramatically reduces false positives from noisy sensor readings.
XI-E Runtime Parameter Tuning via Sliders
While walking, users can fine-tune inference parameters without recompiling:
  • sampleThreshold and windowThreshold sliders adjust the confidence cutoffs.
  • inferenceStride, hitWindow, and requiredHits sliders control how often and how strictly windows are evaluated.
  • minIntervalMs slider sets the minimum gap between detections to prevent chatter.
These sliders live in the "Footfall Tuning" activity, updating the underlying parameters in real time.
XI-F Performance Characteristics
On a BLU JL8 Android device:
  • Inference time: $\sim$7 ms per window.
  • Latency: end-to-ding at 0.5 s.
The android app provides an example for a signal provider that could be consumed by an audio process. The information passed to the audio process would correspond to predetermined points in a song, such as beats. The main, strong footfall signal has been captured here with suitable latency for sending relevant composition information in a timely manner to the audio process.
XII Results and Discussion
As of the date of this release, the mobile app outputs a ding when a footfall is detected, in order to show what’s going on in real time. The footfall signal was successfully picked up with a low enough latency to be usable for music composition. With that said, existing musical protocols such as midi do not suffice to convey the information from the footfall detection control process to the audio playback rendering process.
The audio processing code needs to operate stably in real time, but the control code to pass the predicted footfall information can be more flexible in order to accomodate substantial inference methods like a bi-LSTM. The control code can then pass messages to the C++ audio code without needing to adhere to strict timing constraints. If this approach to predictive music composition turns out to be effective, it might be useful to implement a standard for passing information such as footfall signal between control and rendering process that iterates on Midi 2.0.
One difficulty of this system is that the methods for composing music are very interactive. Digital audio workstations and instruments are rich bidirectional sources of information. In order for this system to be effective, a representation of the internal walking state needs to be clear to the composer so they can design gestures in the music and record multiple parts based on how that signal changes. A better visualization or audiation method borrowed from audio engineering software could be implemented to interact with the footfall signal.
XIII Conclusions and Future Directions
One thing that became clear in doing this research is that music strongly skews towards being a one-way, unilateral form of communication. Specifically, it’s often going to be the most compelling move from an artistic standpoint to capture exactly what the person played in the moment. There are not many existing examples of highly effective, interactive musical compositions.
It’s a worthwhile challenge to create highly interactive music because it hasn’t been done effectively many times. But there’s definitely an increased risk of putting lipstick on a pig, in other words, spending time on a research direction that is unlikely to work.